# PredIN: Towards Open-Set Gesture Recognition via Prediction Inconsistency

### v1

@@ Abstract
Gesture recognition based on surface electromyography (sEMG) has achieved significant progress in human-machine interaction (HMI). However, accurately recognizing predefined gestures within a closed set is still inadequate in practice; a robust open-set system needs to effectively reject unknown gestures while correctly classifying known ones. To handle this challenge, we first report prediction inconsistency discovered for unknown classes due to ensemble diversity, which can significantly facilitate the detection of unknown classes. Based on this insight, we propose an ensemble learning approach, PredIN, to explicitly magnify the prediction inconsistency by enhancing ensemble diversity. Specifically, PredIN maximizes the class feature distribution inconsistency among ensemble members to enhance diversity. Meanwhile, it optimizes inter-class separability within an individual ensemble member to maintain individual performance. Comprehensive experiments on various benchmark datasets demonstrate that the PredIN outperforms state-of-the-art methods by a clear margin. Our proposed method simultaneously achieves accurate closed-set classification for predefined gestures and effective rejection for unknown gestures, exhibiting its efficacy and superiority in open-set gesture recognition based on sEMG.
keywords:
@@
open-set recognition , surface electromyography , ensemble learning , prediction inconsistency , gesture recognition
[…]
1 Introduction
@@
In the human-machine interaction (HMI) paradigm, gesture recognition serves as a foundational task and has been extensively applied across diverse domains [1]. Recently, the development of gesture recognition systems [2, 3, 4] based on surface electromyography (sEMG) signals has been remarkable. However, most of them are confined to closed-set scenarios, where the training and test sets share an identical label space. These closed-set systems lack robustness and reliability in the dynamic and ever-changing real world, which causes them to mistake novel gestures or unintentional motions as known ones and generate false interaction signals. Therefore, a robust gesture recognition system, one which can correctly classify predefined known gestures while identifying unknown gestures in real-world scenarios, is in high demand. Scheirer et al. [5] first described the above demand as open-set recognition (OSR), whose test set contains unknown classes that are not included in the training set.
@@
OSR is an active topic in the field of computer vision, with numerous methods continuously being proposed. However, only a few studies [6, 7] focus on open-set sEMG-based gesture recognition. Due to the inherently random and non-stationary nature of sEMG signals, commonly used methods based on reconstruction or generative models in OSR may not be applicable, particularly in achieving closed-set classification accuracies comparable to discriminative methods [8, 9, 10]. A predominant aspect of existing OSR discriminative methods is to explore the distinctions between known and unknown classes, and design various strategies to enlarge them [11]. Accordingly, a score function is derived based on these distinctions to reject the unknown. A recently popular trend for OSR is employing prototype learning (PL) since it establishes a clear distance distinction between the known and unknown, and demonstrates promising performance [12, 13]. PL methods are able to learn a compact feature space while keeping open space for the unknown.
Beyond the distinction of distance, we reveal that prediction inconsistency within the ensemble learning framework can boost the OSR performance. Within an ensemble learning framework, ensemble members trained with different random initializations can converge to significantly different solutions [14]. This variation causes the ensemble model to perform better than any individual members and more diversity among ensemble members leads to better performance, as they typically do not make the same errors on the same inputs [14], which fulfills one of the OSR tasks’ objectives, a satisfactory closed-set classification ability. Unexpectedly, the ensemble diversity also plays a crucial role in identifying the unknown according to our findings, which accomplishes the other objective of OSR tasks well. In this paper, we first discover prediction inconsistency for unknown samples within the ensemble learning framework. Specifically, ensemble members tend to exhibit inconsistent predictions for the unknown (Fig. 1(b)), while consistently agreeing on the same correct results for known ones (Fig. 1(a)). To better understand the prediction inconsistency, it is important to note that classification models will assign improperly high confidence for unknown samples and misclassify them into known classes [12]. The distinction in prediction inconsistency facilitates the differentiation of unknown samples. In our example of the standard ensemble model, which combines two identical networks, diversity between two members is solely attributed to the randomness of the initialization and the learning procedure [15]. Despite this, it is promising that Fig. 1(c) exhibits pronounced distinctions in prediction inconsistency between known and unknown samples. In light of this, a natural idea is to enhance the ensemble diversity in order to magnify the prediction inconsistency for the unknown.
(a) Known samples
(b) Unknown samples
(c) Prediction inconsistency
@@
Fig. 1: An illustration of prediction inconsistency for the unknown between ensemble members. We summarize the prediction results of samples in the two ensemble members and present them as confusion matrices. Each value in the matrices represents the number of known or unknown samples classified into certain known classes within two members. Both horizontal and vertical coordinates represent class labels. There are pronounced distinctions in prediction inconsistency between known and unknown samples. (a) and (b) represent the prediction results of bioDB2 samples. (c) represents the fraction of prediction inconsistency (%) for both known and unknown samples among four public datasets under the standard ensemble model.
@@
To this end, we propose a new ensemble learning approach, PredIN, to magnify the prediction inconsistency by explicitly enhancing ensemble diversity. Specifically, PredIN introduces two complementary losses which regularize the class feature distribution based on prototype learning [12]. Among ensemble members, PredIN maximizes the inconsistency of class feature distribution by inconsistency loss to enhance diversity. Within an individual ensemble member, PredIN incorporates a triplet loss to optimize inter-class separability, thereby maintaining individual performance. PredIN ultimately rejects the unknown based on prediction inconsistency and distance. We conduct comprehensive experiments on public datasets to validate the superiority of our proposed method. The source code is available at https://github.com/Lchenuu/PredIN.
@@
    We reveal the distinction in prediction inconsistency between known and unknown samples within the ensemble learning framework, which significantly facilitates the rejection of unknown classes.
@@
    Based on the above observation, we propose an ensemble learning approach, PredIN, an effective and explicit diversity-inducing method by maximizing the class feature distribution inconsistency, thereby magnifying the distinction in prediction inconsistency.
@@
    Comprehensive experiments on public sEMG datasets demonstrate that our approach simultaneously maintains closed-set classification accuracy for known gestures and improves rejection for unknown gestures, outperforming previous approaches by a clear margin. Additionally, further experiments on public image datasets demonstrate the general applicability of our approach.
The remainder of this paper is structured as follows. In Section 2, we give the related works based on gesture recognition, open-set recognition and ensemble learning. In Section 3, the details of our proposed PredIN are presented. Extensive evaluations compared with state-of-the-art (SOTA) methods and comprehensive analyses of the proposed approach are reported in Section 4 and Section 5. Finally, the conclusion of this paper is presented in Section 6.
2 Related Works
2.1 Closed-set sEMG-based Gesture Recognition
The emergence of deep learning has freed sEMG-based gesture recognition from the constraints of manual feature extraction [16], facilitating a better understanding of human gestures. Various deep learning architectures have been widely employed for this task. Park et al. [17] pioneered the application of Convolutional Neural Network (CNN) models to classify the Ninapro DB2 dataset [18]. Furthermore, more complex CNN models and Recurrent Neural Network (RNN) models have showcased their superiority to fine gesture classification [2, 3]. EMGHandNet [2] proposed a hybrid CNN and Bi-LSTM framework to capture both the inter-channel and temporal features of sEMG. The attention mechanism is also popular in this field, due to the natural electrode channels and spatial attributes of sEMG signals. Sun et al. [19] proposed a multiscale feature extraction network (MSFENet) based on channel-spatial attention to decode the EMG signals.
@@ 2.2 Open-Set Recognition
Previous OSR discriminative methods established the rejection rules or distinctions mainly in prediction probability [21, 22] and distance [12, 13, 23]. Bendale and Boult [21] demonstrated the limitation of softmax probabilities and introduced OpenMax, a new model prediction layer based on extreme value theory. CPN [12] was the first to introduce prototype learning to OSR, modeling known classes as prototypes and rejecting the unknown based on distance metric. ARPL [13] considered the potential characteristics of the unknown data and proposed the concept of reciprocal points to introduce unknown information. Subsequently, more methods based on prototype learning have been proposed, focusing on improving the compactness of known features [23], mining high-quality and diverse prototypes [24], or constructing multiple Gaussian prototypes for each class [25]. In addition to distance metric, Park et al. [11] observed the distinction in the Jacobian norm between the known and unknown and devised an m-OvR loss to induce strong inter-class separation within the known classes. Numerous researchers believe that modeling only known classes is insufficient and suggest incorporating prior knowledge about unknown classes by generative models. Some approaches attempted to generate fake data [26], counterfactual images [27] or confused samples [13].
@@
Despite advancing OSR performance in image recognition, only a few studies [6, 7] have focused on the challenge of open-set sEMG-based gesture recognition. Wu et al. [6] identified the unknown based on distinctions in distance and reconstruction error through metric learning and autoencoders (AE). To avoid the high computational complexity of generative models, Wu et al. [7] further introduced the convolutional prototype network (CPN) to construct multiple prototypes for known classes, employing a matching-based approach to reject the unknown. While these methods have made progress, there still needs to be further exploration to enhance the performance of open-set sEMG-based gesture recognition.
@@
Different from the above methods, our approach emphasizes the distinction in prediction inconsistency and distance metric to reject the unknown. In light of these distinctions, we propose a discriminative approach based on the ensemble learning framework, which has been rarely explored in OSR tasks. In this paper, we demonstrate that ensemble learning shows superiority not only in known classification but also in unknown rejection. MEDAF [28] also applied an ensemble learning framework to address OSR, but our motivation differs from theirs as they focus on diverse representations, not prediction inconsistency.
2.3 Ensemble Learning
Ensemble methods benefit from the diversity in predictions among ensemble members, as errors made by some members are mitigated by correct predictions from others [14]. Building on this observation, enhancing ensemble diversity has been a persistent focus in ensemble learning. Considering where diversity is injected among ensemble members, we can categorize existing methods into three main types. The first type focuses on the diversity of the input space. They seek to construct different inputs for each member by bagging [29], data augmentation [30], or orthogonal input gradients [31]. The second type works on the diversity of the weight space with the underlying assumption that an ensemble of neural networks with weights distant from each other produces diversified outputs [32]. For example, Repulsive Deep Ensembles [33] introduced a repulsive term to discourage different ensemble members from collapsing to the same function. The third type focuses on the diversity of the output space including the predictions and features. For instance, DICE [15] increased diversity by reducing spurious correlations among features through a mutual information-based method. DBAT [34] enforced the disagreement of predictions on the auxiliary Out of Distribution (OOD) data to promote diversity.
Similarly, our approach explicitly enhances ensemble diversity by maximizing class feature distribution inconsistency in the feature space. Additionally, our method considers maintaining individual performance, an aspect addressed only in DICE. They aim to achieve an optimal balance between ensemble diversity and individual accuracies.
3 Methodology
3.1 Problem Definition
@@
Considering the sEMG-based gesture recognition in real-world scenarios, we assume that $\mathcal{Y}\subset\mathbb{N}$ is the infinite label space of all possible gesture classes. Assume that $\mathcal{C}=\{1,\ldots,N\}\subset\mathcal{Y}$ represents $N$ known classes of interest. The set $\mathcal{U}=\mathcal{Y}\backslash\mathcal{C}$ represents all unknown classes that need to be rejected. The objective of open-set recognition is to find a measurable recognition function $f^{*}\subset\mathbb{H}$ which minimizes both the empirical classification risk on known samples and the open space risk on unknown samples. Open space risk refers to the risk of incorrectly labeling any unknown class as a known one [5].
@@
Fig. 2: An illustration of our proposed framework. Our framework ensembles two members, each of which contains an encoder and a set of learnable prototypes. $\mathcal{L}_{PL}$ and $\mathcal{L}_{trip}$ are applied to each member individually while $\mathcal{L}_{incon}$ simultaneously acts on both. $\mathcal{L}_{incon}$ aims to maximize the class feature distribution inconsistency, ensuring each member has an entirely distinct layout or neighboring class pairs (blue arrows). Upon this, unknown samples like $\mathbf{z}_{u}$ represented by purple are predicted near the clusters of different known classes due to prediction inconsistency while known samples agree on the same predictions across two members.
3.2 Methodology Overview
Combined with prototype learning, we propose an ensemble learning approach, PredIN, to minimize both the empirical classification risk and open space risk simultaneously. Our proposed framework is illustrated in Fig. 2. Specifically, we simultaneously train two ensemble members. In the following text, unless otherwise specified, the number of ensemble members is two. Each member contains an encoder $h$ with an arbitrary architecture and $N$ learnable prototypes of known classes. Prototype learning acts as a basic classification model, with the loss $\mathcal{L}_{PL}$ applied to each member to learn a clear distance distinction. Furthermore, to enlarge the distinction in prediction inconsistency, the inconsistency loss $\mathcal{L}_{incon}$ operates concurrently on both members to enhance their diversity by maximizing the inconsistency of class feature distribution among ensemble members. In addition, we apply the loss $\mathcal{L}_{trip}$ to enhance inter-class separability, thereby maintaining the individual performance.
3.3 Prototype Learning
Each individual member of PredIN employs prototype learning to establish a clear distance distinction. The core idea of PL methods is to encourage samples to be close to their corresponding prototypes and distant from others, where class prototypes serve as centers or representatives of each class [12]. This establishes a compact feature space and a closed classification boundary for known classes while preserving open space for unknown samples [13]. It provides a distance-based approach for rejecting unknown samples, which has been proven superior to softmax-based approaches [24].
For a given sample $\mathbf{x}_{i}$ with label $y_{i}$, its embedding feature is defined as $\mathbf{z}_{i}=h(\mathbf{x}_{i})\in\mathbb{R}^{d}$. $N$ known classes are each assigned a learnable prototype $\mathbf{p}^{k}\in\mathbb{R}^{d}$, where $1\leq k\leq N$. The probability of the prediction result $\hat{y_{i}}$ being $k$ for $\mathbf{x}_{i}$ is based on the distance $d(h(\mathbf{x}_{i}),\mathbf{p}^{k})$:

### v2

# Towards Open-Set Myoelectric Gesture Recognition via Dual-Perspective Inconsistency Learning

@@ Abstract
Gesture recognition based on surface electromyography (sEMG) has achieved significant progress in human-machine interaction (HMI), especially in prosthetic control and movement rehabilitation. However, accurately recognizing predefined gestures within a closed set is still inadequate in practice; a robust open-set system needs to effectively reject unknown gestures while correctly classifying known ones, which is rarely explored in the field of myoelectric gesture recognition. To handle this challenge, we first report a significant distinction in prediction inconsistency discovered for unknown classes, which arises from different perspectives and can substantially enhance open-set recognition performance. Based on this insight, we propose a novel dual-perspective inconsistency learning approach, PredIN, to explicitly magnify the prediction inconsistency by enhancing the inconsistency of class feature distribution within different perspectives. Specifically, PredIN maximizes the class feature distribution inconsistency among the dual perspectives to enhance their differences. Meanwhile, it optimizes inter-class separability within an individual perspective to maintain individual performance. Comprehensive experiments on various benchmark datasets demonstrate that the PredIN outperforms state-of-the-art methods by a clear margin. Our proposed method simultaneously achieves accurate closed-set classification for predefined gestures and effective rejection for unknown gestures, exhibiting its efficacy and superiority in open-set gesture recognition based on sEMG.
Index Terms:
@@
open-set recognition, gesture recognition, dual-perspective, inconsistency learning, surface electromyography.
I Introduction
@@
In the human-machine interaction (HMI) paradigm, gesture recognition serves as a foundational task and has been extensively applied across diverse domains, such as movement rehabilitation [1], prosthetic control [2] and mobile interaction [3]. Recently, the development of gesture recognition systems [4, 5, 6] based on surface electromyography (sEMG) signals has been remarkable. However, most of them are confined to classic closed-set scenarios, where the training and test sets share an identical label space. These closed-set systems lack robustness and reliability in the dynamic and ever-changing real world, which causes them to mistake novel gestures or unintentional motions as known ones and generate false interaction signals. Therefore, a robust gesture recognition system, one which can correctly classify predefined known gestures while identifying unknown gestures in real-world scenarios (Figure 1), is in high demand. Scheirer et al. [7] first described the above demand as open-set recognition (OSR), whose test set contains unknown classes that are not included in the training set.
Figure 1: An illustration of open-set myoelectric gesture recognition system. An open-set myoelectric gesture recognition system need to correctly classify predefined known gestures while rejecting unknown gestures to avoid generating false interaction signals.
@@
OSR is an active topic in the field of computer vision, with numerous methods continuously being proposed. However, only a few studies [8, 9] focus on open-set myoelectric gesture recognition. Due to the inherently random and non-stationary nature of sEMG signals, commonly used methods based on reconstruction or generative models in OSR may not be applicable, particularly in achieving closed-set classification accuracies comparable to discriminative methods [10, 11]. A predominant aspect of existing OSR discriminative methods is to explore the distinctions between known and unknown classes, and design various strategies to enlarge them [12]. Accordingly, a score function is derived based on these distinctions to reject the unknown. A recently popular trend for OSR is employing prototype learning (PL) since it establishes a clear distance distinction between the known and unknown, and demonstrates more promising performance than softmax prediction probabilities distinction [13, 14]. PL methods are able to learn a compact feature space while keeping open space for the unknown. Although the methods based on PL have achieved promising performance, they do not fully explore the inherent distinctions between known and unknown classes as distinguishing them from a single perspective is insufficient.
(a) A toy example
(b) Known samples
(c) Unknown samples
@@
Figure 2: An illustration of prediction inconsistency for the unknown between different perspectives. We summarize the prediction results of samples in the two perspectives and present them as confusion matrices. Each value in the matrices represents the number of known or unknown samples classified into certain known classes within two networks. Both horizontal and vertical coordinates represent class labels. There are pronounced distinctions in prediction inconsistency between known and unknown samples. (a) is a toy example about different perspectives. (b) and (c) represent the prediction results of bioDB2 samples.
Beyond the distinction in distance, we reveal that prediction inconsistency within different perspectives can boost the OSR performance. In real-world scenarios, a single entity can be described in various aspects depending on the perspectives. By considering multiple perspectives, a more reliable assessment of the true similarity between two entities can be made. Unexpectedly, these diverse perspectives play a crucial role in identifying the unknown according to our findings. A toy example is shown in Figure 2(a), we could represent a known sample as a cone and an unknown sample as a cylinder. Imagine projecting the cone and cylinder: the projection of the cylinder appears pseudo-similar to the cone when viewed from the top, but different from another angle. Nonetheless, the projection of the cone retains its resemblance to other cones, regardless of the projection direction. Similarly, a known sample will be consistent with other known samples of the same class from any projection direction, constituting prediction inconsistency (Figure 2(b)). However, it is difficult for an unknown sample to maintain the pseudo-similarity with the same known class across different perspectives, leading to its prediction inconsistency (Figure 2(c)). To better understand the prediction inconsistency, it is important to note that classification models will assign improperly high confidence for unknown samples and misclassify them into known classes [13]. The distinction in prediction inconsistency facilitates the differentiation of unknown samples. In our example, the two perspectives, represented by two identical networks, differ solely due to the randomness of the initialization and the learning procedure [15]. Despite this, it is promising that it exhibits pronounced distinctions in prediction inconsistency between known and unknown samples. Learning to incorporate significantly diverse perspectives will help magnify the prediction inconsistency for the unknown. In light of this, a natural idea is to enhance the differences between perspectives.
@@
To this end, we propose a novel dual-perspective inconsistency learning approach, PredIN, to magnify the prediction inconsistency by explicitly enhancing the inconsistency of class feature distribution within perspectives. Specifically, PredIN constructs two branches representing dual perspectives. Among the two perspectives, PredIN maximizes the inconsistency of class feature distribution by inconsistency loss to enhance perspective differences. Within an individual perspective, PredIN incorporates a triplet loss to optimize inter-class separability, thereby maintaining individual performance. These two strategies act complementarily to regularize the class feature distribution. PredIN ultimately rejects the unknown based on prediction inconsistency and distance. We conduct comprehensive experiments on public datasets to validate the superiority of our proposed method.
@@
    A novel dual-perspective inconsistency learning framework, PredIN, is designed to address a crucial and underexplored task in open-set myoelectric gesture recognition. We reveal a significant distinction between known and unknown samples in open-set scenarios, where prediction inconsistency within different perspectives can notably boost open-set recognition performance.
@@
    To magnify the distinction in prediction inconsistency, we propose two complementary strategies that learn significantly diverse perspectives by explicitly maximizing the class feature distribution inconsistency while maintaining individual performance.
@@
    Comprehensive experiments on multiple public sEMG datasets demonstrate that our approach simultaneously maintains closed-set classification accuracy for known gestures and improves rejection for unknown gestures, outperforming previous approaches by a clear margin.
II Related works
II-A Closed-Set sEMG-based Gesture Recognition
The emergence of deep learning has freed sEMG-based gesture recognition from the constraints of manual feature extraction, facilitating a better understanding of human gestures [16]. Various deep learning architectures have been widely employed for this task. Park et al. [17] pioneered the application of Convolutional Neural Network (CNN) models to classify the Ninapro DB2 dataset [18]. Furthermore, more complex CNN models and Recurrent Neural Network (RNN) models have showcased their superiority to fine gesture classification [16]. EMGHandNet [19] proposed a hybrid CNN and Bi-LSTM framework to capture both the inter-channel and temporal features of sEMG. The attention mechanism is also popular in this field, due to the natural electrode channels and spatial attributes of sEMG signals. Liu et al. [20] proposed a Transformer-based network to capture global contextual information and local details for sparse sEMG recognition.
@@ II-B Open-Set Recognition
Previous OSR discriminative methods established the rejection rules or distinctions mainly in prediction probability [22, 23] and distance [24, 25]. Bendale and Boult [22] demonstrated the limitation of softmax probabilities and introduced OpenMax, a new model prediction layer based on extreme value theory. CPN [13] was the first to introduce prototype learning to OSR, modeling known classes as prototypes and rejecting the unknown based on distance metric. ARPL [14] considered the potential characteristics of the unknown data and proposed the concept of reciprocal points to introduce unknown information. Subsequently, more methods based on prototype learning have been proposed, focusing on improving the compactness of known features [26], mining high-quality and diverse prototypes [27], or constructing multiple Gaussian prototypes for each class [28]. In addition to distance metric, Park et al. [12] observed the distinction in the Jacobian norm between the known and unknown and devised an m-OvR loss to induce strong inter-class separation within the known classes. MEDAF [29] proposed that learning diverse representations is the key factor in OSR. Beyond the discriminative methods, numerous researchers believe that modeling only known classes is insufficient and suggest incorporating prior knowledge about unknown classes by generative models. Some approaches attempted to generate fake data [30], counterfactual images [31] or confused samples [14].
@@
Despite advancing OSR performance in image recognition, only a few studies [8, 9] have focused on the challenge of open-set myoelectric gesture recognition. Wu et al. [8] identified the unknown based on distinctions in distance and reconstruction error through metric learning and autoencoders (AE). To avoid the high computational complexity of generative models, Wu et al. [9] further introduced the convolutional prototype network (CPN) to construct multiple prototypes for known classes, employing a matching-based approach to reject the unknown. While these methods have made progress, there still needs to be further exploration to enhance the performance of open-set sEMG-based gesture recognition.
@@
Different from the above methods, our approach emphasizes the distinction in prediction inconsistency and distance metric to reject the unknown. In light of these distinctions, we propose a discriminative approach based on the dual-perspective inconsistency learning framework, which encourages maximizing differences between perspectives. In this paper, we demonstrate that prediction inconsistency within the dual-perspective inconsistency learning framework shows superiority in unknown rejection.
III METHODOLOGY
III-A Problem Definition
Figure 3: An illustration of our proposed framework. Our framework consists of dual perspectives represented by two branches, each of which contains an encoder and a set of learnable prototypes. $\mathcal{L}_{PL}$ and $\mathcal{L}_{trip}$ are applied to each branch individually while $\mathcal{L}_{incon}$ simultaneously acts on both. $\mathcal{L}_{incon}$ aims to maximize the class feature distribution inconsistency, ensuring each branch has an entirely distinct layout or neighboring class pairs (blue arrows). Upon this, unknown samples like $\mathbf{z}_{u}$ represented by purple are predicted near the clusters of different known classes due to prediction inconsistency while known samples agree on the same predictions across two perspectives.
@@ III-A Problem Definition
Considering the sEMG-based gesture recognition in real-world scenarios, we assume that $\mathcal{Y}\subset\mathbb{N}$ is the infinite label space of all possible gesture classes. Assume that $\mathcal{C}=\{1,\ldots,N\}\subset\mathcal{Y}$ represents $N$ known classes of interest, which appear in both the training and test sets. The set $\mathcal{U}=\mathcal{Y}\backslash\mathcal{C}$ represents all unknown classes that need to be rejected and only appear in the test set. The objective of open-set recognition is to find a measurable recognition function $f^{*}\subset\mathbb{H}$ which minimizes both the empirical classification risk on known samples and the open space risk on unknown samples. Open space risk refers to the risk of incorrectly labeling any unknown class as a known one [7].
@@
III-B Methodology Overview
To minimize both the empirical classification risk and open space risk simultaneously, we propose a novel dual-perspective inconsistency learning approach, PredIN. Our proposed framework is illustrated in Figure 3. Specifically, we simultaneously train two branches representing the two perspectives. Each branch contains an encoder $h$ with an arbitrary architecture and $N$ learnable prototypes of known classes. Prototype learning acts as a basic classification module, with the loss $\mathcal{L}_{PL}$ applied to each branch to learn a clear distance distinction. Apart from serving as a classification module, PL naturally aligns with proposed two strategies of regularizing the class feature distribution (detailed in Section III-D and Section III-E). Specifically, to enlarge the distinction in prediction inconsistency, the inconsistency loss $\mathcal{L}_{incon}$ operates concurrently on both branches to enhance their differences by maximizing the inconsistency of class feature distribution among perspectives. In addition, we apply the loss $\mathcal{L}_{trip}$ to enhance inter-class separability, thereby maintaining the individual performance.
III-C Prototype Learning
Each individual perspective of PredIN employs prototype learning to establish a clear distance distinction. The core idea of PL methods is to encourage samples to be close to their corresponding prototypes and distant from others, where class prototypes serve as centers or representatives of each class [13]. This establishes a compact feature space and a closed classification boundary for known classes while preserving open space for unknown samples [14]. It provides a distance-based approach for rejecting unknown samples, which has been proven superior to softmax-based approaches [27].
For a given sample $\mathbf{x}_{i}\in\mathbb{R}^{C\times T}$ ($C$ denotes the number of sEMG channels and $T$ denotes the number of timesteps) with label $y_{i}$, its embedding feature is defined as $\mathbf{z}_{i}=h(\mathbf{x}_{i})\in\mathbb{R}^{d}$. $N$ known classes are each assigned a learnable prototype $\mathbf{p}^{k}\in\mathbb{R}^{d}$, where $1\leq k\leq N$. The probability of the prediction result $\hat{y_{i}}$ being $k$ for $\mathbf{x}_{i}$ is based on the distance $d(h(\mathbf{x}_{i}),\mathbf{p}^{k})$: